14 April 2015: Tuesday

LDAP:

  • Monday night’s update had a serious problem: OSP sent a birthday that started with a zero (0), which happens to be illegal in the LDAP schema for integers.  The issue was paged at 8:30pm.  I was paged at 9:20pm and worked until 10:30pm to fix the LDAP update issue and run the following processes.
    • Code updated to check for and prevent similar problem in the future.
  • No issues with the ACCOUNT update this morning.
  • Dealing with NetID questions from Account Services.

Backups:

  • Tape and CoM/IS clones finished yesterday.
  • Disk Clones continue with about 17TB left.
  • EMC releases 8.2.1.2 to fix buffer overflow in nsr_render_log command that will allow anyone to run arbitrary commands.
  • Windows client still won’t backup its E drive.  Spent 110 minutes (yes, almost 2 hours) in chat with EMC.  Still not resolved.
  • More work to try to get NFS performance up and going on bujbod2.

 

Tags:

Leave a Reply

You must be logged in to post a comment.